Hi all I am having some problems with my math class I am taking now, but so far I only have one word problem that is giving me grief can someone please help me??
There are 480 dishes that need to be rinsed. John can rinse them in 120 minutes by himself. It will take his friend Bob 80 minutes to rinse these dishes. How long will it take them if they rinse these 480 dishes together?

Answer by Alan3354(21608)

There are 480 dishes that need to be rinsed. John can rinse them in 120 minutes by himself. It will take his friend Bob 80 minutes to rinse these dishes. How long will it take them if they rinse these 480 dishes together?
------------
Find how many dishes they do per minute.
John does 480 dishes/120 minutes = 4/min
Bob does 480/80 = 6/min
Together, they do 10/minute
480/10 = 48 minutes.
---------------
Shortcut = product/sum = 120*80/(120+80) = 9600/200
= 48 minutes